# Can Large Multimodal Models Understand Agricultural Scenes? Benchmarking with AgroMind

### v1

@@ Abstract
Large Multimodal Models (LMMs) has demonstrated capabilities across various domains, but comprehensive benchmarks for agricultural remote sensing (RS) remain scarce. Existing benchmarks designed for agricultural RS scenarios exhibit notable limitations, primarily in terms of insufficient scene diversity in the dataset and oversimplified task design. To bridge this gap, we introduce AgroMind, a comprehensive agricultural remote sensing benchmark covering four task dimensions: spatial perception, object understanding, scene understanding, and scene reasoning, with a total of 13 task types, ranging from crop identification and health monitoring to environmental analysis. We curate a high-quality evaluation set by integrating eight public datasets and one private farmland plot dataset, containing 25,026 QA pairs and 15,556 images. The pipeline begins with multi-source data pre-processing, including collection, format standardization, and annotation refinement. We then generate a diverse set of agriculturally relevant questions through the systematic definition of tasks. Finally, we employ LMMs for inference, generating responses, and performing detailed examinations. We evaluated 18 open-source LMMs and 3 closed-source models on AgroMind. Experiments reveal significant performance gaps, particularly in spatial reasoning and fine-grained recognition, it is notable that human performance lags behind several leading LMMs. By establishing a standardized evaluation framework for agricultural RS, AgroMind reveals the limitations of LMMs in domain knowledge and highlights critical challenges for future work. Data and code can be accessed at https://rssysu.github.io/AgroMind/.
@@ 1 Introduction
Agriculture, as the cornerstone of global food security and economic stability, sustains the livelihoods of the global population and is essential to achieve the Sustainable Development Goal 2 (Zero Hunger) [1, 2]. However, it confronts unprecedented challenges from climate extremes and land degradation. Remote sensing (RS), with its multi-source data integration and real-time monitoring capabilities, provides critical support for crop growth monitoring [3, 4, 5], pest and disease detection [6, 7], and land use change evaluation [8, 9], driving the transformation of agriculture toward intelligence.
Agricultural RS typically involves heterogeneous multi-source data (e.g. optical, SAR, and LiDAR). Traditional deep learning methods struggle with cross-modal fusion, few-shot generalization, and dynamic environmental adaptation [10]. While recent advancements in large multimodal models (LMMs) have significantly enhanced visual understanding and reasoning [11, 12, 13, 14]. As the requirements for visual processing in real-world applications continue to rise, numerous LMMs have been employed to improve the understanding of remote sensing images [15, 16]. The establishment of systematic benchmarking frameworks plays a pivotal role in comprehensively evaluating the core capabilities and latent potential of LMMs.
Despite the impressive capabilities demonstrated by LMMs across various domains, comprehensive benchmarks for agricultural RS remain scarce. Existing benchmarks in the RS field primarily focus on urban scenarios [17, 18, 19], and those designed for agricultural scenarios exhibit notable limitations, primarily in terms of insufficient scene diversity in the dataset and oversimplified task design, as shown in Fig. 2.
@@
To begin with, current benchmarks suffer from insufficient scene diversity in terms of crop types, geographic coverage, and climatic conditions, limiting their ability to reflect the complexity and variability inherent in real-world agricultural environments. Existing datasets in the agriculture field are often limited in scope, primarily focusing on a single crop type or a specific scenario, such as crop diseases [20] or fruit diseases [21], which poses challenges in evaluating model performance in more dynamic and diverse agricultural scenarios (see Table 1). To achieve accurate and comprehensive benchmarking, it is imperative to develop datasets that cover a wide array of crops, diseases, pests, and environmental conditions, addressing the needs of various agricultural applications.
Furthermore, existing benchmarks for agricultural vision tasks often rely on oversimplified task design such as basic classification or detection [20, 22, 23], which fails to capture the complexity of real-world agricultural scenarios. In practice, agricultural applications demand models capable of spatial reasoning, contextual understanding, and support decision-making for tasks like yield forecasting or precision fertilization. While AgroBench[24] introduces a hierarchical task structure that aligns better with practical needs, the lack of quantitative evaluation metrics prevents it from enabling systematic performance assessment of LMMs in agriculture.
@@
RSIEval [25] 10 types 1K
VRSBench [19] 10 types 12K
EarthVQA [26] 6 types 21K
Urbench [17] 14 types 11.6K
LHRS-Bench [27] 11 types >1.1M
XLRS-Bench [18] 16 types 32K
CDDM [20] - 1M
AgriBench [24] 17 types >7K
Agri-LLaVA [22] - 400K
AgroInstruct [28] 6 types 70K
AgMMU [23] 5 types 5K
AgroMind (ours) 13 types 25K
@@
$\bullet$ We construct a large-scale multi-modal and multi-scene agricultural remote sensing dataset, which includes 8 public datasets and 1 private farmland plot dataset, containing 15,556 images and 25,026 QA pairs. The dataset spans critical applications such as pest and disease monitoring, plot analysis, multiple crop recognition, as well as anomaly detection in agricultural scenarios.
@@
$\bullet$ We conduct a comprehensive evaluation of 18 open-source and 3 closed-source LMMs on the proposed AgroMind. Experimental results reveal the limitations of current LMMs in agricultural remote sensing applications, and present the future directions for expert knowledge-based LMMs.
@@ 2.1 Large Multimodal Models
Leveraging the advanced language reasoning and understanding capabilities of Large Language Models (LLMs) [13, 29], LMMs have demonstrated powerful capabilities in multimodal tasks. Recent advancements in LMMs have led to the development of both closed-source systems, such as GPT-4o [11] and Gemini [12], and open-source frameworks like the InternVL series [30], LLaVA series [31, 32], and VILA series [33]. However, despite their success in general tasks, LMMs still face challenges when dealing with domain-specific tasks.
Recent years have seen growing interest in applying LMMs to remote sensing [34, 35], leveraging their ability to integrate visual and textual data for intelligent decision-making. Notable examples include GeoChat [36] and EarthGPT [37], which demonstrate cross-modal capabilities in geospatial analysis. Models such as Agri-LLaVA [22], AgroGPT [28], and WheatRustVQA [38] demonstrate the potential in agricultural tasks like disease diagnosis and crop yield prediction. However, these works primarily focus on narrow scope tasks (e.g., crop classification, disease detection) and often ignore the complexity of agricultural scenarios, such as geospatial reasoning and scene understanding. Our objective is to assess LMMs in a wider variety of agricultural RS tasks, broadening the scope to include more extensive view coverage and fully harnessing their potential in agricultural scenarios.
@@ 2.2 General multimodal benchmark
With the rapid advancement of LMMs, traditional benchmarks such as captioning [39], VQA [40, 41], GQA [42], and TextVQA [43] have become insufficient for evaluating their comprehensive capabilities. Pioneering works such as MME and MMBench established multidimensional evaluation protocols encompassing perception, reasoning, and social cognition through thousands of carefully curated questions. The introduction of MMMU [44] further elevated assessment rigor by incorporating expert-level knowledge evaluation, exposing fundamental limitations in current models’ reasoning capacities. Recently, benchmarks like Seed-Bench [45], MMTBench [46], and MME-Realworld [47] establish large-scale real-world evaluation frameworks. Given the inadequacy of current benchmarks in addressing the unique challenges of specialized fields, there is an urgent need to develop LMM benchmarks that are explicitly tailored to domain-specific knowledge.
@@ 2.3 Remote sensing multimodal benchmark
The RS community has seen a surge in initiatives dedicated to assessing LMMs performance. Multimodal benchmarks have rapidly evolved from early perception-oriented tasks such as image captioning and object classification to more complex reasoning tasks including visual question answering, spatial reasoning, and object relationship understanding. Pioneering works such as EarthVQA [26], RSIEval [25], LHRS-Bench [27], VRSBench [19], UrBench [17], and XLRS-Bench [18] have significantly advanced the evaluation of large multimodal models (LMMs) in general and urban remote sensing scenarios. These efforts have significantly enriched the evaluation of LMMs in general and urban remote sensing contexts.
However, these benchmarks are largely limited to urban, disaster response, or environmental domains, providing insufficient coverage of agricultural remote sensing tasks such as crop identification, growth monitoring, and pest or disease diagnosis. While early efforts like CDDMBench [20] and AgriBench [24] have explored multimodal evaluation in agriculture, they remain constrained by narrow task coverage, limited scene diversity, and small-scale data. Therefore, it is crucial to construct a comprehensive multi-dimensional benchmark for real agricultural RS tasks, to promote the application of LMMs in agricultural intelligence.
@@ 3.1 Dataset statistics
This section presents fundamental statistical characteristics of the AgroMind dataset. The dataset is constructed by integrating eight publicly available RS datasets and one proprietary farmland plot dataset, specifically including OAM-TCD [48], OilPalmUAV [49], AVC [50], PhenoBench [51], Maize Tassel Identification Dataset (Iflytek AI Competition 2024), IP102 [52], ACFR [53], 2018 AI Challenger Dataset, and 7-Provinces Ground Dataset. Through a selection process, we retained 15,556 images and generated 25,026 question-answer (QA) pairs from these visual data, ensuring comprehensive diversity across agricultural scenarios. The proportional contributions from different source datasets are quantitatively illustrated in Fig. 3 (a).
Statistics for different sensors. The AgroMind dataset was compiled from three sensor types: unmanned aerial vehicles (UAVs), satellite platforms, and cameras. Fig. 3 (b) presents the exact numerical distribution of images and corresponding QA pairs for each sensor type, revealing a QA pair ratio of approximately 5:4:1 for ground-based cameras, satellite platforms, and UAVs, respectively. Notably, while satellite and ground imagery constitute the majority of the dataset, UAV-derived annotations still encompass 2,000 QA pairs, ensuring sufficient coverage to validate model robustness in multi-perspective agricultural tasks.
Statistics for different scenes. Fig. 3 (c) demonstrates the dataset’s comprehensive coverage of agricultural scenarios spanning five critical categories: Anomaly Detection, Crop Monitoring, Pest Identification, Plot Delineation, and Tree Analysis. The Crop Monitoring category contains approximately 9,000 high-quality images paired with 4,000 annotated QA pairs, forming the largest data subset. The Tree Analysis and Plot Delineation categories consist of 2,000 images (with 2,000 QA pairs) and 1,000 images (with 3,000 QA pairs), respectively. Notably, the inclusion of specialized scenarios such as Anomaly Detection and Pest Identification demonstrates the dataset’s capacity to address emerging challenges in precision agriculture.
@@ 3.2 Dataset preprocessing
Customized processing protocols were applied to heterogeneous data sources: (1) TIFF-format remote sensing images were converted to PNG/JPG formats to ensure model compatibility and prevent parsing failures; (2) High-resolution geospatial images underwent randomized parcel-boundary cropping to generate multi-scale samples simulating spatial query variations; (3) Instance-level statistical analysis and bounding box annotation were performed using original labels and masks. The resulting standardized dataset contains hierarchical annotations across three levels: pixel-level (segmentation masks), instance-level (object detection boxes), and parcel-level (agricultural boundaries). Original image resolutions were preserved (300×300 to 4,500×4,500 pixels) to maintain data integrity. Complete collection procedures and statistical details are documented in the Appendix B.
@@ 4 AgroMind benchmark
We introduce AgroMind, a comprehensive multimodal benchmark designed to assess the capabilities of LMMs in agricultural scenarios, containing 25,026 image-question-answer pairs. The AgroMind is characterized by the following features: (1) Broad task coverage. AgroMind covers four categories of tasks (Fig. 4(a)): spatial perception, object understanding, scene understanding and reasoning, and a total of 13 task types, covering typical visual language requirements in agricultural production. (2) versatile evaluation. It supports a wide range of question formats such as judgment, counting, image choice, text choice, and open-ended questions, allowing for comprehensive evaluation from basic perception to high-level reasoning (Fig. 4(b)). (3) Domain-specific focus. As visualized in the word clouds, the benchmark emphasizes agricultural-specific terms such as cultivated, land, pest, healthy and anomaly, demonstrating its alignment of the domain (Fig. 4(c)).
@@ 4.2 Benchmark curation
The pipeline of AgroMind covers four key steps from data preparation to quality control, ensuring accurate measurement of LMMs performance in agricultural RS tasks, as detailed in Fig. 6.

### v2

@@ Abstract
Large Multimodal Models (LMMs) has demonstrated capabilities across various domains, but comprehensive benchmarks for agricultural remote sensing (RS) remain scarce. Existing benchmarks designed for agricultural RS scenarios exhibit notable limitations, primarily in terms of insufficient scene diversity in the dataset and oversimplified task design. To bridge this gap, we introduce AgroMind, a comprehensive agricultural remote sensing benchmark covering four task dimensions: spatial perception, object understanding, scene understanding, and scene reasoning, with a total of 13 task types, ranging from crop identification and health monitoring to environmental analysis. We curate a high-quality evaluation set by integrating eight public datasets and one private farmland plot dataset, containing 27,247 QA pairs and 19,615 images. The pipeline begins with multi-source data pre-processing, including collection, format standardization, and annotation refinement. We then generate a diverse set of agriculturally relevant questions through the systematic definition of tasks. Finally, we employ LMMs for inference, generating responses, and performing detailed examinations. We evaluated 20 open-source LMMs and 4 closed-source models on AgroMind. Experiments reveal significant performance gaps, particularly in spatial reasoning and fine-grained recognition, it is notable that human performance lags behind several leading LMMs. By establishing a standardized evaluation framework for agricultural RS, AgroMind reveals the limitations of LMMs in domain knowledge and highlights critical challenges for future work. Data and code can be accessed at https://rssysu.github.io/AgroMind/.
@@ 1 Introduction
Agriculture, as the cornerstone of global food security and economic stability, sustains the livelihoods of the global population and is essential to achieve the Sustainable Development Goal 2 (Zero Hunger) weiss2020remote ; sporchia2024zero . However, it confronts unprecedented challenges from climate extremes and land degradation. Remote sensing (RS), with its multi-source data integration and real-time monitoring capabilities, provides critical support for crop growth monitoring ma2022wheat ; wu2023challenges ; zheng2024review , pest and disease detection wenjiang2019progress ; wang2025advances , and land use change evaluation wang2022evaluation ; yuan2024fusu , driving the transformation of agriculture toward intelligence.
Agricultural RS typically involves heterogeneous multi-source data (e.g. optical, SAR, and LiDAR). Traditional deep learning methods struggle with cross-modal fusion, few-shot generalization, and dynamic environmental adaptation lin2023multimodality . While recent advancements in large multimodal models (LMMs) have significantly enhanced visual understanding and reasoning openai2024gpt4o ; team2023gemini ; touvron2023llama ; bai2023qwen . As the requirements for visual processing in real-world applications continue to rise, numerous LMMs have been employed to improve the understanding of remote sensing images pang2024h2rsvlm ; li2024vision . The establishment of systematic benchmarking frameworks plays a pivotal role in comprehensively evaluating the core capabilities and latent potential of LMMs.
Despite the impressive capabilities demonstrated by LMMs across various domains, comprehensive benchmarks for agricultural RS remain scarce. Existing benchmarks in the RS field primarily focus on urban scenarios zhou2025urbench ; wang2025xlrsbenchmultimodalllmsunderstand ; Li2024VRSBench , and those designed for agricultural scenarios exhibit notable limitations, primarily in terms of insufficient scene diversity in the dataset and oversimplified task design, as shown in Fig. 2.
@@
To begin with, current benchmarks suffer from insufficient scene diversity in terms of crop types, geographic coverage, and climatic conditions, limiting their ability to reflect the complexity and variability inherent in real-world agricultural environments. Existing datasets in the agriculture field are often limited in scope, primarily focusing on a single crop type or a specific scenario, such as crop diseases liu2024cddm or fruit diseases lan2023visual , which poses challenges in evaluating model performance in more dynamic and diverse agricultural scenarios (see Table 1). To achieve accurate and comprehensive benchmarking, it is imperative to develop datasets that cover a wide array of crops, diseases, pests, and environmental conditions, addressing the needs of various agricultural applications.
Furthermore, existing benchmarks for agricultural vision tasks often rely on oversimplified task design such as basic classification or detection liu2024cddm ; wang2024agrillava ; gauba2025agmmu , which fails to capture the complexity of real-world agricultural scenarios. In practice, agricultural applications demand models capable of spatial reasoning, contextual understanding, and support decision-making for tasks like yield forecasting or precision fertilization. While AgroBenchyutong2024agribench introduces a hierarchical task structure that aligns better with practical needs, the lack of quantitative evaluation metrics prevents it from enabling systematic performance assessment of LMMs in agriculture.
@@
RSIEval Hu2025RSGPT 10 types 1K
VRSBench Li2024VRSBench 10 types 12K
EarthVQA Wang2024EarthVQA 6 types 21K
Urbench zhou2025urbench 14 types 11.6K
LHRS-Bench Muhtar2025LHRSBot 11 types >1.1M
XLRS-Bench wang2025xlrsbenchmultimodalllmsunderstand 16 types 32K
CDDM liu2024cddm - 1M
AgriBench yutong2024agribench 17 types >7K
Agri-LLaVA wang2024agrillava - 400K
AgroInstruct Awais2025AgroGPT 6 types 70K
AgMMU gauba2025agmmu 5 types 5K
AgroMind (ours) 13 types 27K
@@
$\bullet$ We construct a large-scale multi-modal and multi-scene agricultural remote sensing dataset, which includes 8 public datasets and 1 private farmland plot dataset, containing 19,615 images and 27,247 QA pairs. The dataset spans critical applications such as pest and disease monitoring, plot analysis, multiple crop recognition, as well as anomaly detection in agricultural scenarios.
@@
$\bullet$ We conduct a comprehensive evaluation of 20 open-source and 4 closed-source LMMs on the proposed AgroMind. Experimental results reveal the limitations of current LMMs in agricultural remote sensing applications, and present the future directions for expert knowledge-based LMMs.
@@ 2.1 Large Multimodal Models
Leveraging the advanced language reasoning and understanding capabilities of Large Language Models (LLMs) touvron2023llama ; achiam2023gpt , LMMs have demonstrated powerful capabilities in multimodal tasks. Recent advancements in LMMs have led to the development of both closed-source systems, such as GPT-4o openai2024gpt4o and Gemini team2023gemini , and open-source frameworks like the InternVL series InternVL2 , LLaVA series LLaVA-NeXT-Interleave ; LLaVA-NeXT , and VILA series lin2024vila . However, despite their success in general tasks, LMMs still face challenges when dealing with domain-specific tasks.
Recent years have seen growing interest in applying LMMs to remote sensing lan2024efficient ; muhtar2024lhrs , leveraging their ability to integrate visual and textual data for intelligent decision-making. Notable examples include GeoChat kuckreja2024geochat and EarthGPT zhang2024earthgpt , which demonstrate cross-modal capabilities in geospatial analysis. Models such as Agri-LLaVA wang2024agrillava , AgroGPT Awais2025AgroGPT , and WheatRustVQA nanavaty2024integrating demonstrate the potential in agricultural tasks like disease diagnosis and crop yield prediction. However, these works primarily focus on narrow scope tasks (e.g., crop classification, disease detection) and often ignore the complexity of agricultural scenarios, such as geospatial reasoning and scene understanding. Our objective is to assess LMMs in a wider variety of agricultural RS tasks, broadening the scope to include more extensive view coverage and fully harnessing their potential in agricultural scenarios.
@@ 2.2 General multimodal benchmark
With the rapid advancement of LMMs, traditional benchmarks such as captioning agrawal2019nocaps , VQA marino2019ok ; goyal2017making , GQA hudson2019gqa , and TextVQA yang2021tap have become insufficient for evaluating their comprehensive capabilities. Pioneering works such as MME and MMBench established multidimensional evaluation protocols encompassing perception, reasoning, and social cognition through thousands of carefully curated questions. The introduction of MMMU yue2024mmmu further elevated assessment rigor by incorporating expert-level knowledge evaluation, exposing fundamental limitations in current models’ reasoning capacities. Recently, benchmarks like Seed-Bench li2024seed , MMTBench ying2024mmt , and MME-Realworld zhang2024mme establish large-scale real-world evaluation frameworks. Given the inadequacy of current benchmarks in addressing the unique challenges of specialized fields, there is an urgent need to develop LMM benchmarks that are explicitly tailored to domain-specific knowledge.
@@ 2.3 Remote sensing multimodal benchmark
The RS community has seen a surge in initiatives dedicated to assessing LMMs performance. Multimodal benchmarks have rapidly evolved from early perception-oriented tasks such as image captioning and object classification to more complex reasoning tasks including visual question answering, spatial reasoning, and object relationship understanding. Pioneering works such as EarthVQA Wang2024EarthVQA , RSIEval Hu2025RSGPT , LHRS-Bench Muhtar2025LHRSBot , VRSBench Li2024VRSBench , UrBench zhou2025urbench , and XLRS-Bench wang2025xlrsbenchmultimodalllmsunderstand have significantly advanced the evaluation of large multimodal models (LMMs) in general and urban remote sensing scenarios. These efforts have significantly enriched the evaluation of LMMs in general and urban remote sensing contexts.
However, these benchmarks are largely limited to urban, disaster response, or environmental domains, providing insufficient coverage of agricultural remote sensing tasks such as crop identification, growth monitoring, and pest or disease diagnosis. While early efforts like CDDMBench liu2024cddm and AgriBench yutong2024agribench have explored multimodal evaluation in agriculture, they remain constrained by narrow task coverage, limited scene diversity, and small-scale data. Therefore, it is crucial to construct a comprehensive multi-dimensional benchmark for real agricultural RS tasks, to promote the application of LMMs in agricultural intelligence.
@@ 3.1 Dataset statistics
This section presents fundamental statistical characteristics of the AgroMind dataset. The dataset is constructed by integrating eight publicly available RS datasets and one proprietary farmland plot dataset, specifically including OAM-TCD veitchmichaelis2024oamtcdgloballydiversedataset , OilPalmUAV zheng2021growing , AVC Chiu_2020_CVPR , PhenoBench weyler2024pami , Maize Tassel Identification Dataset (Iflytek AI Competition 2024), IP102 Wu2019Insect , ACFR bargoti2016deep , 2018 AI Challenger Dataset, and 7-Provinces Ground Dataset. Through a selection process, we retained 19,615 images and generated 27,247 question-answer (QA) pairs from these visual data, ensuring comprehensive diversity across agricultural scenarios. The proportional contributions from different source datasets are quantitatively illustrated in Fig. 3 (a). More details can be found in Appendix B.1.
Statistics for different sensors. The AgroMind dataset was compiled from three sensor types: unmanned aerial vehicles (UAVs), satellite platforms, and cameras. Fig. 3 (b) presents the exact numerical distribution of images and corresponding QA pairs for each sensor type, revealing a QA pair ratio of approximately 9:11:7 for ground-based cameras, satellite platforms, and UAVs, respectively. Notably, while satellite and ground imagery constitute the majority of the dataset, UAV-derived annotations still encompass 7,000 QA pairs, ensuring sufficient coverage to validate model robustness in multi-perspective agricultural tasks.
Statistics for different scenes. Fig. 3 (c) demonstrates the dataset’s comprehensive coverage of agricultural scenarios spanning five critical categories: Anomaly Detection, Crop Monitoring, Pest Identification, Plot Delineation, and Tree Analysis. The Crop Monitoring category contains approximately 2,000 high-quality images paired with 7,000 annotated QA pairs, forming the largest data subset. The Pest Identification and Tree Analysis categories consist of 5,000 images (with 6,000 QA pairs) and 6,000 images (with 6,000 QA pairs), respectively. Notably, the inclusion of specialized scenarios such as Anomaly Detection and Plot Delineation demonstrates the dataset’s capacity to address emerging challenges in precision agriculture.
@@ 3.2 Dataset preprocessing
Customized processing protocols were applied to heterogeneous data sources: (1) TIFF-format remote sensing images were converted to PNG/JPG formats to ensure model compatibility and prevent parsing failures; (2) High-resolution geospatial images underwent randomized parcel-boundary cropping to generate multi-scale samples simulating spatial query variations; (3) A manual screening step was performed on augmented images to remove samples exhibiting artifacts such as content overlap or excessive flipping that could distort semantic information;(4) Instance-level statistical analysis and bounding box annotation were performed using original labels and masks. The resulting standardized dataset contains hierarchical annotations across three levels: pixel-level (segmentation masks), instance-level (object detection boxes), and parcel-level (agricultural boundaries). Original image resolutions were preserved (300×300 to 4,500×4,500 pixels) to maintain data integrity. Moreover, we applied a logic enhancement strategy in the QA pair design, artificially predefining the hierarchical reasoning process instead of directly using the raw annotations as the answers of QA pairs. Complete collection procedures and statistical details are documented in the Appendix B.
@@ 4 AgroMind benchmark
We introduce AgroMind, a comprehensive multimodal benchmark designed to assess the capabilities of LMMs in agricultural scenarios, containing 27,247 image-question-answer pairs. The AgroMind is characterized by the following features: (1) Broad task coverage. AgroMind covers four categories of tasks (Fig. 4(a)): spatial perception, object understanding, scene understanding and reasoning, and a total of 13 task types, covering typical visual language requirements in agricultural production. (2) versatile evaluation. It supports a wide range of question formats such as judgment, counting, image choice, text choice, and open-ended questions, allowing for comprehensive evaluation from basic perception to high-level reasoning (Fig. 4(b)). (3) Domain-specific focus. As visualized in the word clouds, the benchmark emphasizes agricultural-specific terms such as cultivated, land, pest, healthy and anomaly, demonstrating its alignment of the domain (Fig. 4(c)).
@@ 4.2 Benchmark curation
The pipeline of AgroMind covers four key steps from data preparation to quality control, ensuring accurate measurement of LMMs performance in agricultural RS tasks, as detailed in Fig. 6. More details can be found in Appendix B and C.